suppose that two pipes (A and B) can be used to fill a swimming pool. If both pipes are used together then it takes 25 hours to fill the pool. If only pipe A is used then it takes 40 hours to fill the pool. If only Pipe B is used the it takes x hours to fill the pool. What formula can be used to find the answer?
Answer by mananth(16949)

Filling the pool is 1 job
Pipe A + pipe B takes 25 hours
So they will do 1/25of the job in 1 hour
Pipe A alone takes 40 hours
It does 1/40 of the job in 1 hour
If Pipe A does the job in x hours it does 1/x of the job in 1hour
1/25-1/40=1/x
LCD = 200
(8-5)/200=1/x
3/200=1/x
Cross multiply
3x=200
x=67 hours